9. A certain electromagnetic wave has a maximum electric field of 1,000 newtons per coulomb. If the intensity of this wave is 1.34 x 103 watts, calculate the maximum magnetic field. If this light in incident directly on a perfect reflector with an area of one square meter for a period of 10.0 seconds, what is the momentum provided to the surface?
Answer #1

a) I = [Emax * Bmax]/2\muo

so; Bmax = 2I\muo /Emax = [2 * 4\pi*10-7 * 1.34*103]/1000 = 3.36\muT

b) for a perfect reflector; momentum, p =2U/c

where U is the total energy =u/\Delta T

where u = B2/2\muo = (3.36*10-6)2/(2*4*3.14*10-7) = 4.49*10-6J

and U = 4.49*10-6/10 = 4.49*10-7J/s

and p = 4.49*10-7 / 3*108 = 1.49*10-15 J/m